OpenStudy (richyw):

When we say a particle has a certain energy. Is this the total energy, or just the kinetic energy? This still confuses me. Basically what I want to know is how to find the momentum of a 3MeV electron!

OpenStudy (anonymous):

this will be total energy.... 3Mev is total energy .E=p*c..where E is total energy p is momentum and c is speed of light..

OpenStudy (anonymous):

find P(momentum).

OpenStudy (richyw):

Still confused because if it is total energy then why wouldn't I have to subtract the rest mass from that?

OpenStudy (fwizbang):

Usually, if it just says energy, you should assume its the total energy. For an electron, which has a mass, the total energy and momentum are related by E^2 = (pc)^2 +(mc^2)^2

OpenStudy (richyw):

ok I have the KINETIC ENERGY of a particle. what is its MOMENTUM?!?! clearly I don't get it at all. I see those formulas everywhere. the kinetic energy is 0.025, it's a nuetron. what is its momentum.

OpenStudy (richyw):

sorry I am just really frustrated with switching between relativaistic and non relatavistic and all this BS. I just need to know how to plug numbers into the relatavistic formula

OpenStudy (fwizbang):

OK. What are the units on the 0.025?

OpenStudy (richyw):

eV. I know I don't need to use the relatavistic definition, but I would rather anyways all the time or I get confused!

OpenStudy (richyw):

for example in another question the energy of an electron is 3.5 TeV so I will need the formula for that one too!

OpenStudy (fwizbang):

So, for this problem you don't need the relativistic formula. The way to tell is to compare the kinetic energy of the particle to the rest energy(mc^2). For a neutron, the rest energy is 940 Mega-eV(MeV), which is a billion times bigger than .025 eV. When the rest energy is really big compared to the KE, the particle is moving very slowly compared to the speed of light, and you can get away with the non relativistic expressions for the KE, p^2/2m. So p = root(2M E) .

OpenStudy (fwizbang):

For the electron, mc^2= .511 MeV, which is a lot smaller than 3.5 TeV, so you need to use the full relativistic expression That put up originally and solve for p.

OpenStudy (richyw):

alright thanks a lot. That is what I cannot do haha!

OpenStudy (fwizbang):

E^2 =(Pc)^2 +(mc^2)^2 (Pc)^2=E^2-(mc^2)^2 Pc= root(E^2-(mc^2)^2) P= (root(E^2-(mc^2)^2))/c

OpenStudy (richyw):

ok cool so that is where I got. now my question (actualyl original question is what is E? is that the kinetic energy?

OpenStudy (richyw):

because in one question I have been given the kinetic energy of an electron. In the other question they just was "3Mv electrons".

OpenStudy (richyw):

also thanks so much medal coming your way for sure for letting me know i'm at least on the right track

OpenStudy (fwizbang):

Unfortunately, the answer depends on the context(which sucks), but if they don't specify the kinetic energy explicitly, you should probably assume its the total energy. Sometimes there are clues, like in the case of the neutron, where 0.025 eV was much less than the neutron's energy, so the 0.025 eV had to be KE, because E>=mc^2 always.

OpenStudy (richyw):

alright thanks a lot. It's too late for this assignment but I will try to figure it out after class!